### v1

# eprint Nisho-2-2025

Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current primarily dissipates the axion energy at the surface. In contrast, if the cylinder is composed of a material with low conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the axion energy is dissipated mainly inside the bulk of the cylinder. Within the QCD axion model, the dissipated power $P$ is estimated as $P\simeq 2.8\times 10^{-28}\text{W}g_{\gamma}^{2}\left(L/100\text{cm}\right)\left(R/2\text{cm}\right)^{2}\left(m_{a}/10^{-4}\text{eV}\right)\left(B/10\text{T}\right)^{2}\left(10/\epsilon\right)\left(\rho_{a}/0.3\text{GeVcm}^{-3}\right)$, with $\sigma=10^{-3}$eV, radius $R$, length $L$, electric permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. Using an LC circuit appropriately tuned to a quality factor $Q=10^{6}$, the signal-to-noise ratio at low temperature 100100mK and observation time 6060sec. is so large $\sim 10$ that the detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
A central issue in particle physics is to identify phenomena beyond the Standard Model. The axion, proposed as the Nambu–Goldstone boson of Peccei–Quinn symmetry axion ; axion1 ; axion2 , provides a natural solution to the strong CP problem and is also a well-motivated dark matter candidate. It is called as QCD axion. The viable mass window for the QCD axion is tightly constrained to $m_{a}=10^{-6}\mbox{-}10^{-3}\text{eV}$ Wil ; Wil1 ; Wil2 .
Numerous experiments are underway to search for dark matter axions review , most exploiting axion–photon conversion in a strong magnetic field. The induced electromagnetic radiation is expected to be detected with resonant cavitiesadmx ; haystac , superconducting qubitsmoroi , Quantum Hall effect iwazaki ; aiwazaki , e.t.c.
In this letter we propose a new method for axion detection using a cylindrical sample with low electric conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
Dark matter axion generates an oscillating electric field in the presence of a strong magnetic field, which in turn induces an oscillating current in the cylinder. By applying the magnetic field parallel to the cylinder axis, the induced current flows parallel to the external field.
In general, such currents are confined within the skin depth $\delta$. For highly conductive materials, e.g. $\sigma=10^{4}\text{eV}$, the skin depth is extremely small ( $\delta\sim 10^{-5}\text{cm}$ for $m_{a}=10^{-4}$eV ), restricting the current to a thin surface layer. As a result, the effective volume occupied by current flow is small, and the dissipated power from Joule heating is very low to be difficult to detect.
On the other hand, for a cylinder with low conductivity the induced current flows the bulk, and the resulting dissipated power $P$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ). We show that
$P\simeq 2.8\times 10^{-27}\mbox{W}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\text{eV}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
with $\sigma=10m_{a}=10^{-2}$eV, where length L=100L=100cm and radius R=2R=2cm of the cylinder with electric permittivity $\epsilon=10$. The power $P$ depends slightly on $\sigma$ so that even with $\sigma=5\times 10^{-3}$eV, the power is still large $P\simeq 2.2\times 10^{-27}W$. Uniquely, the axion–electromagnetic coupling drives an oscillating current that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents are confined to the surface at high frequencies.
By adopting an appropriate conductivity, e.g. $\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 100$mK and radius ( length ) of the cylinder 22cm ( 100100cm ), we show that using LC circuit tuned with a quality factor $Q=10^{6}$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with large signal-noise ratio $\sim 10$.
Our method using cylinder with low electric conductivity is also effective for so called, dark photon. The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the axion that oscillating electric current is induced in the whole of cylinder.
The main difficulty in detecting axion dark matter lies in its extremely weak coupling to electromagnetic radiation or ordinary matter ( electrons or nucleons ). In particular, the interaction between the axion field $a(t,\vec{x})$ and the electromagnetic field is described by
$L_{a\gamma\gamma}=g_{a\gamma\gamma}a(t,\vec{x})\vec{E}\cdot\vec{B},$ (2)
where $\vec{E}$ and $\vec{B}$ denote the electric and magnetic fields, respectively. The coupling constant is $g_{a\gamma\gamma}=g_{\gamma}\alpha/\pi f_{a}$, with the fine structure constant $\alpha\simeq 1/137$, the axion decay constant $f_{a}$, and the relation $m_{a}f_{a}\simeq 6\times 10^{-6}\text{eV}\times 10^{12}\text{GeV}$ for the QCD axion model. The model dependent coefficient is $g_{\gamma}\simeq 0.37$ for the DFSZ model dfsz ; dfsz1 and $g_{\gamma}\simeq-0.96$ for the KSVZ model ksvz ; ksvz1 .
For a classical axion field representing dark matter, the interaction term $g_{a\gamma\gamma}a(t,\vec{x})$ is extremely small. Assuming that dark matter consists entirely of axions, the local energy density of the dark matter axion is
$\rho_{d}=m_{a}^{2}\overline{a(t,\vec{x})^{2}}=\frac{1}{2}m_{a}^{2}a_{0}^{2}\simeq 0.3\text{GeV}/\text{cm}^{3},$ (3)
where the overline denotes time averaging. This yields an effective CP-violating interaction of order $g_{a\gamma\gamma}a(t,\vec{x})\sim 10^{-21}$, essentially independent of the QCD axion mass. Consequently, the axion-induced electric field in vacuum under an external magnetic field $B_{0}$ is extremely weak, of order $\sim g_{a\gamma\gamma}aB_{0}$.
In this letter we show that, for a cylindrical sample in Fig.1 with low electrical conductivity, the Joule heating induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$. In this case, the dissipation occurs throughout the bulk of the cylinder, while for a good conductor with large conductivity, the dissipation is confined to a thin surface layer.
We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 10/m_{a}$, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$. In particular, we fix the radius R=2R=2cm thought out the paper as a practical value. As we show soon later, the condition of $\sigma m_{a}=\epsilon$ maximize the power dissipated inside the cylinder with electric permittivity $\epsilon$. Such a matter with low conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with suitable impurity doping.
A strong external magnetic field $\vec{B}_{0}$ is applied parallel to the cylinder axis, so that the system is axially symmetric. $\vec{B}_{0}=(0,0,B_{0})$ in cylindrical coordinates $(\rho,\theta,z)$, with $\rho=0$ at the center and $\rho=R$ at the surface of the cylinder. In the presence of dark matter axion, the magnetic field induces an oscillating current parallel to $\vec{B}_{0}$. This current produces microwave radiation with frequencies corresponding to axion masses in the range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
We calculate the axion-induced electric field to obtain oscillation electric current and its dissipation power $P$.
Figure 1: cylinder sample with length $L$ and radius $R$ under external magnetic field $\vec{B}_{0}$
In order to obtain oscillating electric current $\vec{J}=\sigma\vec{E}^{\prime}$, we solve the Maxwell equations involving axion effect,
$\vec{\partial}\cdot(\epsilon\vec{E}^{\prime}+g_{a\gamma\gamma}a(t,\vec{x})\vec{B})$ $=0$ $,\quad\vec{\partial}\times\Big(\vec{B}-g_{a\gamma\gamma}a(t,\vec{x})\vec{E}^{\prime}\Big)-\partial_{t}\Big(\epsilon\vec{E}^{\prime}+g_{a\gamma\gamma}a(t,\vec{x})\vec{B}\Big)=\vec{J},$ (4)
$\vec{\partial}\cdot\vec{B}$ $=0$ $,\quad\vec{\partial}\times\vec{E}^{\prime}+\partial_{t}\vec{B}=0$ (5)
where electric permittivity $\epsilon$ and electric current $\vec{J}$ are non vanishing inside the cylinder, while $\epsilon=1$ and $\vec{J}=0$ in vacuum. We have assumed trivial permeability $\mu=1$ of the cylinder.
The electric field $\vec{E}^{\prime}$ induced by the axion effect is much small, being of the order of $g_{a\gamma\gamma}aB_{0}$, while $\vec{B}=\vec{B}_{0}+\vec{B}^{\prime}$ with external magnetic field $\vec{B}_{0}=(0,0,B_{0})$. $\vec{B}^{\prime}$ is induced by the axion and associated with $\vec{E}^{\prime}$.
It is easy to obtain the following equation of electric field $\vec{E}^{\prime}$ using cylindrical coordinate.
$\big(\partial_{\rho}^{2}+\frac{1}{\rho}\partial_{\rho}+\epsilon m_{a}^{2}+i\sigma m_{a}\big)\vec{E}^{\prime}=-m_{a}^{2}g_{a\gamma\gamma}a(t)\vec{B}_{0}$ (6)
assuming $a(t)\propto\exp(-im_{a}t)$. The solution $\vec{E}^{\prime}=(0,0,E^{\prime})$ is
$E^{\prime}=d(t)J_{0}(bm_{a}\rho)+\frac{m_{a}^{2}E_{a}}{\epsilon m_{a}^{2}+im_{a}\sigma}$ (7)
with $E_{a}\equiv-g_{a\gamma\gamma}a(t)B_{0}$, where $d(t)\propto\exp(-im_{a}T)$ is a constant determined by boundary conditions at $\rho=R$. The corresponding magnetic field $\vec{B}^{\prime}=(0,B^{\prime},0)$ is given by solving $\partial_{t}B^{\prime}=\partial_{\rho}E^{\prime}$,
$B^{\prime}=-ib\,d(t)J_{1}(bm_{a}\rho)$ (8)
In the above expression, $J_{0,1}(x)$ denotes Bessel function of the first kind and is chosen because of its finiteness of $E^{\prime}(\rho)$ at $\rho=0$. The constant $b$ is given by
$b\equiv(\epsilon^{2}+y^{2})^{1/4}\exp(i\theta/2)\quad\mbox{with}\quad\theta=\cos^{-1}\Big(\frac{\epsilon}{\sqrt{\epsilon^{2}+y^{2}}}\Big)$ (9)
with $y\equiv\sigma/m_{a}$.
These solutions $E^{\prime}$ and $B^{\prime}$ represent electric and magnetic fields inside the cylinder with radius $R$. Similarly, we can find solutions $E_{v}$ and $B_{v}$ outside the cylinder with the conditions $\vec{J}=0$ and $\epsilon=1$ in the Maxwell equations.
$E_{v}=\tilde{d}(t)H_{0}^{(1)}(m_{a}\rho)+E_{a}\quad\mbox{and}\quad B_{v}=-i\tilde{d}(t)H_{1}^{(1)}(m_{a}\rho)$ (10)
with $H_{0,1}^{(1)}$ Hankel functions of the first kind, where $\tilde{d}\propto\exp(-im_{a}t)$ is a constant determined by boundary conditions at $\rho=R$. The Hankel function of the first kind is chosen because the radiations described by $E_{v}$ and $B_{v}$ are outgoing waves; $E_{v}(B_{v})\sim\exp(-im_{a}t+im_{a}\rho)$ as $\rho\to\infty$.
In order to determine the constants $d(t)$ and $\tilde{d}(t)$, we impose boundary conditions at the surface $\rho=R$ of the cylinder such that $\epsilon E^{\prime}(\rho=R)=E_{v}(\rho=R)$ and $B^{\prime}(\rho=R)=B_{v}(\rho=R)$. Then, we have
$\epsilon d(t)J_{0}(bm_{a}R)+\frac{\epsilon m_{a}^{2}E_{a}}{\epsilon m_{a}^{2}+im_{a}\sigma}=\tilde{d}(t)H_{0}^{(1)}(m_{a}R)+E_{a}\quad\mbox{and}\quad b\,d(t)J_{1}(bm_{a}R)=\tilde{d}(t)H_{1}^{(1)}(m_{a}R)$ (11)
Therefore, by solving the equations for $d(t)$ and $\tilde{d}(t)$, we have
$d(t)=\frac{iy}{\epsilon+iy}\frac{E_{a}H_{1}^{(1)}(x)}{\epsilon J_{0}(bx)H_{1}^{(1)}(x)-bJ_{1}(bx)H_{0}^{(1)}(x)},\quad\tilde{d}(t)=d(t)\Bigg(\frac{bJ_{1}(bx)}{H_{1}^{(1)}(x)}\Bigg),$ (12)
with $x=m_{a}R$.
Thus, the oscillating electric and magnetic fields $E^{\prime}$ and $B^{\prime}$ inside the cylinder are
$E^{\prime}(\rho)$ $=$ $E_{a}\Big(\frac{1}{\epsilon+iy}+\frac{iy}{\epsilon+iy}\frac{H_{1}^{(1)}(m_{a}R)J_{0}(bm_{a}\rho)}{\epsilon J_{0}(bm_{a}R)H_{1}^{(1)}(m_{a}R)-bJ_{1}(bm_{a}R)H_{0}^{(1)}(m_{a}R)}\Big)$ (13)
$B^{\prime}(\rho)$ $=$ $E_{a}\Big(\frac{by}{\epsilon+iy}\frac{E_{a}H_{1}^{(1)}(m_{a}R)J_{1}(bm_{a}\rho)}{\epsilon J_{0}(bm_{a}R)H_{1}^{(1)}(m_{a}R)-bJ_{1}(bm_{a}R)H_{0}^{(1)}(m_{a}R)}\Big).$ (14)
They oscillate as $E^{\prime}(B^{\prime})\propto\exp(-im_{a}t)$. It is easy to see that the first term of the electric field, $E^{\prime}(\rho)$, exists throughout the bulk of the cylinder, whereas the second term is confined to its surface; $J_{0}(bm_{a}\rho)/J_{0,1}(bm_{a}R)\propto\exp\big((R-\rho)m_{a}\text{Im}(b)\big)$ where $\text{Im}(O)$ denotes imaginary part of $O$.
Uniquely, the axion–electromagnetic coupling drives an oscillating electric field that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where the electric field is confined to the surface at high frequencies. A similar fact is also present in the case of dark photonkishimoto .
As shown later, the second term dominates the power dissipation in the limit of high conductivity $(\sigma\sim 10^{4}\text{eV})$, while the first term becomes dominant for low conductivity $(\sigma\sim 10^{-3}\text{eV})$.
Using the formula of the electric field $E^{\prime}$, we obtain oscillating electric current $J(\rho)=\sigma Re(E^{\prime}(\rho))$ and corresponding its power $P$ averaged over the period $2\pi/m_{a}$,
$P=\int_{0}^{R}\overline{J(\rho)(Re(E^{\prime}(\rho)))}L2\pi\rho d\rho=\int_{0}^{R}\sigma|E^{\prime}(\rho)|^{2}L\pi\rho d\rho$ (15)
where $Re(O)$ denotes real part of the quantity $O$.
When we put $z=\rho/R$, the formula $P$ is rewritten such that
$P=\sigma L\pi\int_{0}^{1}|E^{\prime}(Rz)|^{2}R^{2}zdz=\frac{\pi L|E_{a}|^{2}}{m_{a}}\int_{0}^{1}yx^{2}|U(x,y,z)|^{2}zdz,$ (16)
where $U(x,y,z)$ is
$U(x,y,z)\equiv\frac{1}{\epsilon+iy}+\frac{iy}{\epsilon+iy}\frac{H_{1}^{(1)}(x)J_{0}(bxz)}{\epsilon J_{0}(bx)H_{1}^{(1)}(x)-bJ_{1}(bx)H_{0}^{(1)}(x)}$ (17)
We remind that $y\equiv\sigma/m_{a}$ and $x\equiv m_{a}R$ and $b=b(y)$ is the function of $y$. $P$ is a complicated function in $x$ and $y$, or $m_{a}$ and $R$. But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small $\sigma\sim 10^{-3}$eV.
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}m_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with depth $\delta$ of the cylinder. It is coincide with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
The skin depth for general $\sigma$ is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$. In this paper, assuming typical value $\epsilon=10$ of semiconductor, we take low electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) at temperature $\sim 100$mK and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$. The matter may be fabricated of for instance a semiconductor with appropriate doping of impurities.
In such a matter with low electric conductivity, we find that the main contribution to $P$ comes from the first term in $U$. Actually, the following quantities are monotonically increasing function in $x$,
$\int_{0}^{1}$ $yx^{2}dzz\Bigg|\frac{iy}{\epsilon+iy}\frac{H_{1}^{(1)}(x)J_{0}(bxz)}{\epsilon J_{0}(bx)H_{1}^{(1)}(x)-bJ_{1}(bx)H_{0}^{(1)}(x)}\Bigg|^{2}=0.01\text{-}0.09\,\,\,\mbox{for}\,\,\,x=1\text{-}10$ (18)
$\int_{0}^{1}$ $yx^{2}dzz\Big|\frac{1}{\epsilon+iy}\Big|^{2}=0.025\text{-}2.5\,\,\,\mbox{for}\,\,\,x=1\text{-}10.$ (19)
with $y=10$ and $\epsilon=10$. We find that for each $x$, the quantity in the second equation is larger than the one in the first equation. Especially, it is more than an order of magnitude for $x\geq 10$. Further, smaller $y$ ( <10<10 ) leads to much larger discrepancy between the quantities. Therefore, we may write
$P=\frac{\pi L|E_{a}|^{2}}{m_{a}}\int_{0}^{1}yx^{2}|U(x,y,z)|^{2}zdz\sim\frac{\pi L|E_{a}|^{2}}{2m_{a}}\Big(\frac{yx^{2}}{\epsilon^{2}+y^{2}}\Big).$ (20)
$P$ takes the maximal value at $y=\sigma/m_{a}=\epsilon$ and $P(\epsilon,y=\epsilon)\propto\epsilon^{-1}$. Thus, it is favorable to take the conductivity $\sigma=\epsilon m_{a}$. Because we do not know the value of the axion mass $m_{a}$, the value $\sigma=\epsilon m_{a}$ is unknown. But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$. In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, when we use semiconductor with $\epsilon\sim 10$, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV. Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection. Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
Numerically, we find
$P(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-28}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (21)
with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 10$ ( R=2R=2cm ). Obviously, $P$ in eq.(20) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross-sectional area $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
For comparison, we present the power $P$ when the conductivity is much high, $\sigma=10^{4}$eV,
$P(\sigma=10^{4}\text{eV})$ $\simeq$ $|E_{a}|^{2}\frac{\pi m_{a}\delta RL}{2}\Bigg|\frac{H_{1}^{(1)}(x)}{H_{0}^{(1)}(x)}\Bigg|^{2}$ (22)
$\simeq$ $8.0\times 10^{-32}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (23)
with $y=\sigma/m_{a}=10^{8}$ and $x=m_{a}R\simeq 10$. It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder. We find that the power $P(\sigma=10^{-3}\text{eV})$ with low conductivity, is three order of magnitude larger than $P_{h}(\sigma=10^{4}\text{eV})$ with high conductivity. That is, $P(\sigma=10^{-3}\text{eV})>10^{3}P(\sigma=10^{4}\text{eV})$. It is the reason why we use such a cylinder with low conductivity $\sigma\sim 10^{-3}$eV.
We compare $P$ with thermal noise $P_{t}=T\delta\omega/2\pi$ with the width $\delta\omega=10^{-6}m_{a}$ of the microwave frequency,
$P_{t}\simeq 3.3\times 10^{-20}W\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)$ (24)
We have
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}\simeq 8.5\times 10^{-9}g_{\gamma}^{2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (25)
We use LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$ to detect the power $P$ induced by the axion. When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}^{-1}\sqrt{L_{c}/C_{c}}=10^{6}$, the signal to noise ( S/N ) ratio is given by
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 10\,g_{\gamma}^{2}\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}}$ (26)
with $y=10$ and $x\simeq 10$, where $t_{ob}$ denotes the observation time.
When we change electric conductivity $\sigma=10^{-3}$eV slightly such as $\sigma=5\times 10^{-4}$eV, the S/N ratio $\simeq 8.2$. The dependence on the conductivity is very weak. But, when axion mass is smaller such as $m_{a}=10^{-5}$eV, it is favorable to use smaller $\sigma=10^{-4}$eV to maximize $P$. Then, it leads to
$\frac{P(\sigma=10^{-4}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 3.2g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (27)
with $y=10$ and $x\simeq 1$.
Furthermore, when axion mass is larger such as $m_{a}=10^{-3}$eV, it is favorable to use $\sigma=10^{-2}$eV to maximize $P$. It leads to
$\frac{P(\sigma=10^{-2}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 32\,g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-3}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (28)
with $y=10$ and $x\simeq 100$. Even when $\sigma=5\times 10^{-3}$eV, the S/N ratio is $\simeq 26$. We find that the dependence on the conductivity is also weak.
It turns out that we can search the region of the axion mass $m_{a}=10^{-4}\text{-}10^{-3}$eV with large S/N ratio. In particular, even by fixing $\sigma=5\times 10^{-3}$eV and R=2R=2cm, the S/N ratios are $\simeq 8.2$ for $m_{a}=10^{-4}$eV and $\simeq 26$ for $m_{a}=10^{-3}$eV. The dependence of S/N ratio on the conductivity $\sigma$ is weak so that even if the conductivity of the sample is not uniform in the large sample ( L=1L=1m and R=2R=2cm ), the signal of the axion dark matter could be found in the method.
Obviously, the method of the axion detection using cylinder sample with small electric conductivity is very powerful. This is because the axion–electromagnetic coupling induces an oscillating electric current throughout the entire cylindrical sample, producing sufficiently large dissipation power $P$ for detection.
Summarized, we have proposed a method for axion detection using a cylindrical sample with choosing appropriate low conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$. For such conductivity, the axion-induced electric current penetrates the bulk, yielding substantial energy dissipation and a high signal-to-noise ratio, demonstrating the feasibility of detecting dark matter axions under realistic experimental conditions. The method is also effective for the search of dark photon.